OpenStudy (cenaida):

A rectangle measures 96 inches long and 72 inches wide. What is the diagonal length of the rectangle? A. 125.1 in B. 120 in C. 106.2 in D. 10 in

OpenStudy (amistre64):

how do we find the lengths associated with a right trianlge?

OpenStudy (cenaida):

I don't know..

OpenStudy (amistre64):

what does your material suggest?

OpenStudy (cenaida):

do I multiply 96 and 72?

OpenStudy (amistre64):

not quite, im trying to get something more pythagorean from this

OpenStudy (amistre64):



OpenStudy (cenaida):

a2+b2+c2?

OpenStudy (cenaida):

@amistre64

OpenStudy (amistre64):

assuming a typo, yes: a^2 + b^2 = c^2 so c = sqrt(a^2 + b^2)

OpenStudy (amistre64):

what is the value of sqrt(96^2 + 72^2)

OpenStudy (cenaida):

B. 120in^2 :)